OpenStudy (anonymous):

I need to solve for 0<=x<=2pi 8sin^2x+6sinx=9

OpenStudy (anonymous):

If you let y=sinx, then that's like solving this equation: 8y^2 + 6y = 9

OpenStudy (anonymous):

The solutions you get for this equation will be in terms of "y" so don't forget to substitute y=sinx and you'll have solve those equations.

OpenStudy (anonymous):

I got the factors and put sin back in for x. Now I am stuck again. What do I do with the first factor sin x= 3/4?

OpenStudy (anonymous):

Take the inverse sine of 3/4 : x = arcsin(3/4) is the exact answer.

OpenStudy (anonymous):

Ohhhhhh! I remember that now. geez thanks.

OpenStudy (anonymous):

Thanks for helping me.

OpenStudy (anonymous):

you're welcome....

OpenStudy (anonymous):

So is the arcs in for this = pi/3 just to be sure I got it? ( as you can tell I am a bit lost)